OpenStudy (anonymous):

Given the system of equations presented here: 4x + y = 4 2x + 7y = 28 Which of the following actions creates an equivalent system such that, when combined with the other equation, one of the variables is eliminated? Multiply the second equation by −1 to get −2x − 7y = −28 Multiply the second equation by −4 to get −8x − 28y = −112 Multiply the first equation by −7 to get −28x − 7y = −28 Multiply the first equation by −2 to get −8x − 2y = −8

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

I really don't know, that's why I need help

OpenStudy (anonymous):

Okedoke you are looking for the answer that will change one equation in a way that makes either x y or = cancel out with the other equation. for example if one equation is: 5x + 7y = 28 and the second one is -5x +6y = 100 the variable that cancels out is X. because -5 and 5

OpenStudy (anonymous):

Okay, so would it be D because if it is multiplied by -2 the x would cancel out?

OpenStudy (anonymous):

so in your case the only equation that would make either x,y or = cancel out is the third one because: -28x -7y = -28 2x + 7y =28 see how the y variable cancelled?

OpenStudy (anonymous):

Oh okay, thank you

OpenStudy (anonymous):

it actually cancels y and the =